# General task optimal planning for heterogeneous teams with precedence and compatibility constraints and its application on power grid inspection using unmanned aerial vehicles

### v1

@@ 2.2 Mixed-integer programming problem (MILP) representation
It forces that $p_{u|w}\leq p_{v|w}$ unless ${y_{u|w}=0}$ or ${y_{u|w}=0}$, i.e, either $u$ or $v$ is not visited by $w$ for which case the constraint becomes trivial.
Consider that instead of order constraints imposed directly into specific vertices we need to implement then into task visits in any of the approaches. This is done by imposing Equation 18 in all the pairs of tasked vertices generated by pairing approaches from the first task and from the second one. This can be done for all workers if the limitation is not specific to a worker.
At this point, the model provides no information about the local, per step, status of the workers. This means that although the total cost of a route is known, no information about the partial sum of all the previous visits up to a specific vertex is known. Inspired by [11], some new variables and constraints that keep track of this information are proposed for any monotonically increasing (or decreasing) partial cost function $f^{(\mu)}_{w}:V\,\rightarrow\mathbb{R}$ ($\mu$ is any of the cost types, $w\in W$). The following constraints ($\mathcal{C}_{\text{MFE}}$) are valid and fully determine ${\mathbf{f}=\{f^{(\mu)}_{w}(v)\in\mathbb{R}\}_{w\in W;\,v\in V}}$:
@@
and the three auxiliary constraints:
@@
where $O_{w}^{(\mu)}$ is any prefixed value that fulfils ${O_{w}^{(\mu)}>\sum_{e\in E\lvert_{w}}\Omega^{(\mu)}_{e}}$ (larger than any possible possible route for $w$).
Note that these are a generalization to the MTZ SECs Equation 15, Equation 16 and Equation 17. This means that they are directly usable with the DFJ SECs and that they by themselves can be also used as SECs if the correct cost type is selected. They also assign $f^{(\mu)}_{w}(u)=0$ if $u$ is not visited by $w$.
As an example, consider the time cost type $\mu=t$. As each value of the function $f^{(t)}_{w}(v)$ is encoding the time at which each visited vertex is completed and is therefore monotonically increasing, the previous constraints can be used to obtain its values internally within the solving process and without any external manipulation to the MILP.
With these values available to the MILP solver, constraints can also be imposed locally. Consider the aforementioned case where a vertex must be completed before another one is visited. We can now impose this constraint based on timing instead of route order. We denote this condition as $u\preceq v$, i.e. the time at which $u$ is completed is equal or previous to the time at which $v$ is visited, no matter which worker completes any of them (or $u\prec v$ for the strict inequality). This can be imposed using one precedence constraint ($\mathcal{C}_{\text{P}}$) for each precedence-constrained pair $u,v\in V$:
@@
These type of parameters can also be used as formulated to keep track of the partial energy costs (or equivalently, used or remaining battery charge).
With all of this in mind, we define the space of all possible choices of the values for the free parameters ${\mathbf{\Phi}=\mathbf{Z}\cup\mathbf{Y}\cup\mathbf{P}\cup\mathbf{f}}$ that fulfil all of this constraints as the space of feasible solutions $\mathcal{S}_{F}$. With a description for what a feasible solution is, we need to define a metric that selects the optimal solution out of the entire space $\mathcal{S}_{F}$. Such metric is given as a function ${F:\mathcal{S}_{F}\rightarrow\mathbb{R}}$ that maps each feasible solution $\mathbf{\Phi}\in\mathcal{S}_{F}$ to a real value that needs to be minimal (for maximization, we just change the sign of the cost function and minimize the resulting function).
The specific expression for the cost function heavily depends on the application. However, we give an example that cover the vast majority of cases.
@@
This function however is non-linear and can not be included in MILPs. To deal with this, we a linearization of the function that preserves its characterization. We introduce a new free real parameter $M_{\Sigma}$. We then set $M_{\Sigma}$ as the upper bound for each of the routes times and minimize its value. Instead of directly minimizing the slowest work time, we so do indirectly via this new parameter. All of this can be done by adding a new set of constraints ($\mathcal{C}_{\text{MTM}}$):
@@
This MILP can be solved with any of the off-the-shelf solvers such as CMPLX or SCIP[12] or with any of the standard well-established algorithms. We warn however that this formulation is ultimately based on the Travelling Salesman Problem which is NP-complete [13], although no proof is given, it is assumable that this problem is also NP-complete. Some optimizations will be needed in order to solve the problem in a competent time and more so in real time applications. One possible optimization is the aforementioned relaxation of the subtour constraints using the DFJ SECs.
@@ 2.3 Example I: Classical TSPs
In Table 1, the solving statistics of several TSPLIB problems are shown. The instance file were directly downloaded from [1] and modified to suit our solver format. In general, only one base is used and it is assigned to the point in the original file. The original distance scaling was maintain and the speed at which workers move is set to $10$ units of distance per second in all cases. All the problems are solved as a regular TSP, with only one worker and no time cost assigned to the visit process. Then, some new instances are created with new constraints and modifications. In some cases, the problem is solved with several workers, with compatibility constraints and with order/precedence constraints. Each instance has been solved $10$ times using both the DJF and the MTZ and then the average solving time for each case has been computed. In the former case, the dynamic relax-solve-iterate solving approach has been use for subroute subsets with cardinality less than $8$ to avoid extremely long iteration processes. All the variables and constraints related to order and precedence are also removed. For the MTZ case, the complete model is used. All MILPs has been solved using the default parameters and methods of the SCIP 8.0.4 Optimization Suite.
@@
Table 1: Solving statistics for several instances of classical TSPLIB problems. $n_{w}$ denotes the number of workers, $n_{v}$ the number of vertices, $n_{\text{cc}}$ the number of compatibility constraints, $n_{\text{oc}}$ the number of order constraints,$n_{\text{pc}}$ the number of precedence constraints, $\langle t_{\text{DFJ}}\rangle$ and $\langle t_{\text{MTZ}}\rangle$ are the average compute time of 10 runs using either the DFJ or the MTZ SECs. “Cost” indicates the value of the cost function for the optimal solution. ”D.N.F.” stands for Did Not Finish before the $10000$ s time limit while “-” indicates that such case can not be solved or is not implemented in our code.
@@
The cases where just one worker is active and no additional constraint is added reduce to an asymmetric version of the original TSP, i.e., a TSP where $2$ $z$ variables per edge are used. The run time for such cases are one or two order of magnitude slower that the fastest regular TSP exact algorithms [14] due to the duplicate variables and the lack of optimization.
@@
In general, the use of the DFJ SECs significantly reduces the computation time. This is due to lack of the $\mathcal{P}$ variables. However, in some cases the use of MTZ is actually better. Consider the case with $n_{w}=3$ for the eil22 and eil23. As the number of worker increases, so does the probability of finding subtours in the relaxed problem. Although no proof is given, this can be seen just by doing simple combinatorics on the original TSP version of the problems and checking how the space of feasible solution increases with $n_{w}$ while keeping track of the fraction of solution with no subtours. As such fractions decreases, so does the chance of reaching a subtour free solution of the relaxed problem in the dynamic DFJ solution process. At some point, the number of relaxed and semirelaxed instances needed to completely solve the problem is large enough to surpass the time needed to solve the complete problem with MTZ SECs. This argument is even stronger when considering that the number of possible DFJ constraints increases exponentially while the number of MTZ SECs increases as a polynomial.
@@
It can also be seen that the variable that most affects the complexity of the problem is the number of workers of the instance. Just by going from $1$ to $2$ the compute time jumps several order of magnitude. In many TSP-based models that consider each worker as a unique entity such as the HMWTPP, the number of variables increases linearly with the number of workers. As the number of potential solutions increases exponentially (as $2^{n}$ for problems with only binary variables) with the number of variables, the computation time also potentially scales up. The addition of an additional worker, does not increase the solution time as much, however, it results in several instances not being solved before the $10000$ s time limit.
@@
On the other hand, the compatibility constrains cut the solving time at least by an order of magnitude as it simplifies the problem by eliminating many feasible solutions from original problem. The results for the order and precedence constrains are mixed. In general, the precedence constraints are harder to solve as the parameters involved are real numbers instead of binary parameters. In both cases however, the problem is solved faster as the number of constrains increases. As it happened with the compatibility constrains, this is due to smaller feasible solution space although instead of having a simpler graph and MILP, it is the pre-processing done by the SCIP solver that eliminates the invalid solutions.
@@
In any case, the main objective of this paper is not numerical performance nor analysis. Several papers already do this in much detail [14, 2, 3, 15]. Out point in this section is just to estimate the level of performance of our naive approach to solving the proposed HMWTPP. The computational efficiency obtained in this paper still need to be improved with state-of-the-art algorithms and a future paper is proposed to address this. It might also be interesting to analyze the performance of non-exact and heursitcs algorithms with HMWTPP instances.
@@ 2.4 Example II: Guitar assembly line with 2 workers
To show the versatility of our formulation we will solve a trivial example that models an automated assembly line but includes many of the unique characteristics specific to our model. Consider the manufacturing of a guitar where three task-specific workbenches, WB1111, WB2222 and WB3333, are used by two different workers, $w_{a}$ and $w_{b}$, to complete the steps needed to fully assemble a guitar. The list of tasks, compatibility and time cost for both workers are:
@@
    Time Costs $\rightarrow$ $\{w_{b}:\text{ N.C.},\,w_{a}:\,2\text{ h}\}$
@@
    Time Costs $\rightarrow$ $\{w_{b}:\text{ N.C.},\,w_{a}:\,2\text{ h}\}$
@@
    Time Costs App. A: $\rightarrow$ $\{w_{a}:\,0.5\text{ h},\,w_{b}:\,1\text{ h}\}$
    Time Costs App. B: $\rightarrow$ $\{w_{a}:\,3\text{ h},\,w_{b}:\,3\text{ h}\}$
@@
    Time Costs $\rightarrow$ $\{w_{a}:\,1\text{ h},\,w_{b}:\,2\text{ h}\}$
@@
The solution given by solving the corresponding MILP [^3] All the computations in this paper were made using custom Python and SCIP code available at [16] in a single core of an AMD Ryzen R5555 $5600$X at stock clocks and $32$ Gb of RAM at $3200$ MHz. using the MTM cost function Equation 26 is represented at 1(a) with the slowest route being $7$ hours and $10$ seconds with a computation time of just $50$ ms. This is however not the actual fastest solution.
@@
Due to how the MILP problem is formulated, each workers must leave each task immediately after completion. This means that workers can not wait for others to complete their tasks. Ideally, the faster worker should wait for the slower at T2222 and then complete T5555 and T6666. Instead, the slowest one is who completes the final tasks yielding a slower than optimal solution.
@@
To fix this, the addition of waiting points in proposed. A waiting point is a new abstract non-mandatory task whose completion time cost $\Omega^{\text{WP}}_{w}\geq 0$ (one for each worker $w\in W$) is variable and controlled by the MILP. When this or similar cases are given to the solver, the new variable cost allows the workers to wait a certain amount of time to resume their route, alleviating the aforementioned issue. To keep things simple, these waiting points will be merged into the bases, so no new vertices are added and the only modification needed to be done is to add $\Omega^{\text{WP}}_{w}$ to all the edges that come out of the base. As the cost functions do not explicitly depend on when each task is completed, the waiting process can be done at any point in the route so the value of $\Omega^{\text{WP}}_{w}$ represents the total wait time of the route. As such, we just need to modify the constrains Equation 21 and Equation 22 by doing the following transformation: ${f^{(t)}_{w}(v)\rightarrow f^{(t)}_{w}(v)-\Omega^{\text{WP}}_{w}}$ which just translates the time origin for each UAV independently. We also need ensure that the selected value of ${O_{w}^{(t)}}$ is enough to cover such transformation. When and how much time to wait depends almost exclusively on the user solving the problem with the only restrictions being that the precedence constraints must be fulfilled and the total wait time is equal to the value given by the solver.
@@
The solution with this addition is represented at 1(b) and yields a final time of $5.5$ hours and $5$ seconds with a wait time of $1$ hour, $59$ minutes and $55$ seconds, more than $1.5$ hours faster than the previous solution with a computation time of just $200$ ms.
@@
The basic HMWTPP is a very flexible and robust framework to model many different planification problems. It is clear however that the specific adaptation to certain problems may require some kind of extension of the original model. We believe however that by using the HMWTPP, this process is much direct and less complex than the one required by using some other model. With this example, we show that in can be used in a generic assembly line but many other applications exits with none or trivial modifications. Consider for example the scheduling of human resources in a project management setting, the distribution of data within a network, autonomous storage or even in autonomous shipment delivery. In the last two cases, the constraints Equation 19, Equation 20, Equation 21 and Equation 22 would need to be extended to any type of partial cost function, monotonic or not. Then, they can be used to keep track of the physical weight of the packages delivered or moved by the workers and their battery charge. Simultaneously, the HMWTPP already allows the use of time-windows constraints. Consider that we need a package to be delivered to a certain shop in between to fixed times, otherwise, the delivery is not possible. This is easily done using a constraint of the type:

### v2

@@ 2.2 Mixed-integer programming problem (MILP) representation
It forces that $p_{u|w}\leq p_{v|w}$ unless ${y_{u|w}=0}$ or ${y_{u|w}=0}$, i.e., either $u$ or $v$ is not visited by $w$ for which case the constraint becomes trivial.
Consider that instead of order constraints imposed directly into specific vertices we need to implement them into task visits in any of the approaches. This is done by imposing Equation 18 in all the pairs of tasked vertices generated by pairing approaches from the first task and from the second one. This can be done for all workers if the limitation is not specific to a worker.
At this point, the model provides no information about the local, per step, status of the workers. This means that although the total cost of a route is known, no information about the partial sum of all the previous visits up to a specific vertex is known. Inspired by [9], some new variables and constraints that keep track of this information are proposed for any monotonically increasing (or decreasing by changing the inequalities) partial cost function $f^{(\mu)}_{w}:V\,\rightarrow\mathbb{R}$ ($\mu$ is any of the cost types, $w\in W$). The following constraints ($\mathcal{C}_{\text{MFE}}$) are valid and fully determine ${\mathbf{f}=\{f^{(\mu)}_{w}(v)\in\mathbb{R}\}_{w\in W;\,v\in V}}$:
@@
and the three auxiliary constraint types:
@@
where $O_{w}^{(\mu)}$ is any prefixed value that fulfills ${O_{w}^{(\mu)}>\sum_{e\in E\lvert_{w}}\Omega^{(\mu)}_{e}}$ (it is larger than any possible complete route cost for $w$).
Note that these are a generalization of the MTZ SECs Equation 15, Equation 16 and Equation 17. This means that they by themselves can be also used as SECs if the correct cost type is selected. They also assign $f^{(\mu)}_{w}(u)=0$ if $u$ is not visited by $w$.
As an example, consider the time cost type $\mu=t$. As each value of the function $f^{(t)}_{w}(v)$ is encoding the time at which each visited vertex is completed and is therefore monotonically increasing, the previous constraints can be used to obtain its values internally within the solving process and without any external manipulation after the solution is found.
With these values available to the MILP solver, constraints can also be imposed locally as a per step/vertex basis. Consider the aforementioned case where a vertex must be completed before another one is visited. We can now impose this constraint based on timing instead of route order. We denote this condition as $u\preceq v$, i.e. the time at which $u$ is completed is equal or previous to the time at which $v$ is firstly visited, no matter which worker completes any of them (or $u\prec v$ for the strict inequality). This can be imposed using one precedence constraint ($\mathcal{C}_{\text{P}}$) for each precedence-constrained pair $[u,v]\in V$:
@@
This type of parameters can also be used as formulated to keep track of the partial energy costs (or equivalently, used or remaining battery charge).
With all of this in mind, we define the space of all possible choices of the values for the free parameters ${\mathbf{\Phi}=\mathbf{Z}\cup\mathbf{Y}\cup\mathbf{P}\cup\mathbf{f}}$ that fulfill all of these constraints as the space of feasible solutions $\mathcal{S}_{F}$. With a description of what a feasible solution is, we need to define a metric that selects the optimal solution out of the entire space $\mathcal{S}_{F}$. Such metric must be given as a function ${F:\mathcal{S}_{F}\rightarrow\mathbb{R}}$ that maps each feasible solution $\mathbf{\Phi}\in\mathcal{S}_{F}$ to a real value that needs to be minimal (for maximization, we just change the sign of the cost function and minimize the resulting function).
The specific expression for the cost function heavily depends on the application. However, we give an example that covers many cases.
@@
This function however is non-linear and cannot be included in MILPs. To deal with this, we define a linearization of the function that preserves its characterization. We introduce a new free real parameter $M_{\Sigma}$. We then set $M_{\Sigma}$ as the upper bound for each of the routes’ times and minimize its value. Instead of directly minimizing the slowest work time, we so do indirectly via this new parameter. All of this can be done by adding a new set of constraints ($\mathcal{C}_{\text{MTM}}$):
@@
This MILP can be solved with any of the off-the-shelf solvers such as CMPLX or SCIP[16] or with any of the standard well-established algorithms. We warn however that, as already mentioned, this formulation can be assumed to be at least NP-hard. Some optimizations will be needed in order to solve the problem in a competent time and more so in real-time applications. One possible optimization is the aforementioned dynamical relaxation of the subtour constraints using the DFJ SECs.
@@ 2.3 Example I: Classical TSPs
In Table 1, the solving statistics of several TSPLIB problems are shown. The instance files were directly downloaded from [12] and modified to suit our solver format. In general, only one base is used and it is assigned to the first point in the original file. The original distance scaling was maintained and the speed at which workers move is set to $10$ units of distance per second in all cases. All the problems are solved as a regular TSP, with no time cost assigned to the visit process. Then, some new instances are created with new constraints and modifications that are either randomly generated or by hand by grouping vertices by their number within the instance. In some cases, the problem is solved with several workers, with compatibility constraints and with order/precedence constraints. Each instance has been solved $10$ times using both the DJF and the MTZ and then the average solving time for each case has been computed. In the former case, the dynamic relax-solve-iterate solving approach has been used for subroute subsets with cardinality less than $8$ to avoid too long iteration processes. All the variables and constraints related to order and precedence are also removed in those cases. For the MTZ case, the complete model is used. All MILPs have been solved using the default parameters and methods of the SCIP 8.0.4 Optimization Suite. All the computations in this paper were made using custom Python and SCIP code and problem instances available at [17] in a single core of an AMD Ryzen R5555 $5600$X at stock clocks and $32$ Gb of DDR4 RAM at $3200$ MHz.
@@
Table 1: Solving statistics for several instances of classical TSPLIB problems. $n_{w}$ denotes the number of workers, $n_{v}$ the number of vertices, $n_{\text{cc}}$ the number of compatibility constraints, $n_{\text{oc}}$ the number of order constraints,$n_{\text{pc}}$ the number of precedence constraints, $\langle t_{\text{DFJ}}\rangle$ and $\langle t_{\text{MTZ}}\rangle$ are the average compute time of 10 runs using either the DFJ or the MTZ SECs. “Cost” indicates the value of the cost function for the optimal solution. ”D.N.F.” stands for Did Not Finish before the $10000$ s time limit while “-” indicates that such a case cannot be solved or is not implemented in our code.
@@
The cases where just one worker is active and no additional constraint is added reduce to an asymmetric version of the original TSP, i.e., a TSP where two $z$ variables per edge are used. The run times for such cases are one or two order of magnitude slower that the fastest regular TSP exact algorithms [18] due to the duplicate variables and the lack of optimization.
@@
In general, it can be seen that the use of the DFJ SECs significantly reduces the computation time up to a certain number of vertices. This is due to lack of the $\mathbf{P}$ variables. However, in some cases the use of MTZ is actually better. Consider the case with $n_{w}=3$ for the eil22 and eil23. As the number of workers increases, so does the probability of finding subtours in the relaxed problem. Although no proof is given, this can be seen just by doing simple combinatorics on the original TSP version of the problems and checking how the space of feasible solution increases with $n_{w}$ while keeping track of the fraction of solution with no subtours. As such fraction decreases, so does the chance of reaching a subtour free solution of the relaxed problem in the dynamic DFJ solution process. At some point, the number of relaxed and semi-relaxed instances needed to completely solve the problem is large enough to surpass the time needed to solve the complete problem with MTZ SECs. This is even stronger when considering that the number of possible DFJ constraints increases exponentially while the number of MTZ SECs increases as a polynomial.
@@
It can also be seen that the variable that most affects the complexity of the problem is the number of workers of the instance. Just by going from $1$ to $2$ the compute time jumps several orders of magnitude. In many TSP-based models that consider each worker as a unique entity such as the HMWTPP, the number of variables increases linearly with the number of workers. As the number of potential solutions increases exponentially (as $2^{n}$ for problems with only binary variables) with the number of variables, the computation time also scales up on average. The addition of another worker might result in several instances not being solved before the $10000$ s time limit.
@@
On the other hand, the compatibility constraints cut the solving time at least by an order of magnitude as it simplifies the problem by eliminating many feasible solutions from original problem. The results for the order and precedence constraints are mixed however. In general, the precedence constraints are harder to solve as the parameters involved are real numbers instead of integer parameters. In both cases however, the problem is solved faster as the number of constrains increases. As it happened with the compatibility constraints, this is due to smaller feasible solution space although instead of having a simpler graph and MILP, it is the pre-processing done by the SCIP solver that eliminates the invalid solutions.
Figure 1: Solution for the eil51 instance with two workers and 50 compatibility constraints
@@
In any case, the main objective of this paper is not the numerical performance nor solution analysis. Several papers already do this in much detail [18, 2, 3, 19]. Our point in this section is just to estimate the level of performance of our naive approach to solving the proposed HMWTPP. The computational efficiency obtained in this paper still needs to be improved with state-of-the-art algorithms and a future paper is proposed to address this. It might also be interesting to analyze the performance of non-exact and heuristics algorithms with HMWTPP instances.
@@ 2.4 Example II: Guitar assembly line with 2 workers
To show the versatility of our formulation we will solve a trivial example that models an automated assembly line but includes many of the unique characteristics specific to our model. Consider the manufacturing of a guitar where three task-specific workbenches, WB1111, WB2222 and WB3333 (whose distance is negligible), are used by two different workers, $w_{a}$ and $w_{b}$, to complete the steps needed to fully assemble a guitar. The list of tasks, compatibility and time cost for both workers are:
@@
    $\Rightarrow$ Time Costs $\rightarrow$ $\{w_{b}:\text{ N.C.},\,w_{a}:\,2\text{ h}\}$
@@
    $\Rightarrow$ Time Costs $\rightarrow$ $\{w_{b}:\text{ N.C.},\,w_{a}:\,2\text{ h}\}$
@@
    $\Rightarrow$ Time Costs App. A: $\rightarrow$ $\{w_{a}:\,0.5\text{ h},\,w_{b}:\,1\text{ h}\}$
    $\Rightarrow$ Time Costs App. B: $\rightarrow$ $\{w_{a}:\,3\text{ h},\,w_{b}:\,3\text{ h}\}$
@@
    $\Rightarrow$ Time Costs $\rightarrow$ $\{w_{a}:\,1\text{ h},\,w_{b}:\,2\text{ h}\}$
@@
The solution given by solving the corresponding MILP using the MTM cost function Equation 26 is represented at 2(a) with the slowest route being $7$ hours and $10$ seconds with a computation time of just $50$ ms. This is however not the actual fastest solution.
@@
Due to how the MILP problem is formulated, each worker must leave each task immediately after completion. This means that workers cannot wait for others to complete their tasks. Ideally, the faster worker should wait for the slower at T2222 and then complete T5555 and T6666. Instead, the slowest one is the one who completes the final tasks yielding a slower than optimal solution.
@@
To fix this, the addition of waiting points in proposed. A waiting point is a new abstract non-mandatory task whose completion time cost $\Omega^{\text{WP}}_{w}\geq 0$ (one for each worker $w\in W$) is variable and controlled by the MILP. When this or any similar cases are given to the solver, the new variable cost allows the workers to wait a certain amount of time to resume their route, alleviating the aforementioned issue. To keep things simple, these waiting points will be merged into the bases, so no new vertices are added and the only modification needed to be done is to add $\Omega^{\text{WP}}_{w}$ to all the edges time cost that come out of the base. As the cost function does not explicitly depend on when each task is completed, the waiting process can be done at any point in the route so the value of $\Omega^{\text{WP}}_{w}$ actually represents the total wait time of the route that that later be distributed between several vertices with enough care. As such, we just need to modify the constrains Equation 21 and Equation 22 by doing the following transformation: ${f^{(t)}_{w}(v)\rightarrow f^{(t)}_{w}(v)-\Omega^{\text{WP}}_{w}}$ which translates the time origin for each UAV independently. We also need ensure that the selected value of ${O_{w}^{(t)}}$ is enough to cover such transformation. When and how much time to wait depends almost exclusively on the user solving the problem with the only restrictions being that the precedence constraints must be fulfilled and the total wait time must be equal to the value given by the solver.
@@
The solution with this addition is represented at 2(b) and yields a final time of $5.5$ hours and $5$ seconds with a wait time of $1$ hour, $59$ minutes and $55$ seconds, more than $1.5$ hours faster than the previous solution with a computation time of $200$ ms.
@@
The basic HMWTPP is a very flexible and robust framework to model many different planning problems. It is clear however that the specific adaptation to certain problems may require some kind of extension of the original model. We believe however that by using the HMWTPP, this process is much more direct and less complex than the one that would be required by using some other model. With this example, we show that it can be used in a generic assembly line but many other applications exist with none or trivial modifications. Consider for example the scheduling of human resources in a project management setting, the distribution of data within a network, autonomous storage or even autonomous shipment delivery. In the last two cases, the constraints Equation 19, Equation 20, Equation 21 and Equation 22 would need to be extended to any type of partial cost function, monotonic or not. Then, they can be used to keep track of the physical weight of the packages delivered or moved by the workers and their battery charge. Simultaneously, the HMWTPP already allows the use of time-windows constraints. Consider that we need a package to be delivered to a certain shop within a time windows defined by two fixed time values, otherwise, the delivery is not possible. This is easily done using a constraint of the type: